Find the inverse of 77 modulo 5 by using extended Euclidean Algorithm



Step by step solution


1
Expert's answer
2021-11-05T10:32:35-0400

Apply Euclidean Algorithm to compute GCD as shown in the left column.

It will verify that "GCD(77, 5)=1."

Then we will solve for the remainders in the right column.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 77=5(15)+2 & 2=77-5(15) \\ 
 5=2(2)+1 & 1=5-2(2) \\
 2=2(1)+0 & 
\end{array}"

Use the equations in the right side and perform reverse operation as:


"\begin{matrix}
 1=5-2(2)\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ \\
 \ \ \ 1=5-[77-5(15)](2)\\ \\
1=5(31)+77(-2)\ \ \ 
\end{matrix}"

Therefore "1=77(-2)\mod {5}," or of we prefer a residue value for multiplicative inverse


"1=77(3)\mod {5}."

Therefore, "3" is the multiplicative inverse of "77."